# LiRA: A Multi-Agent Framework for Reliable and Readable Literature Review Generation

### v1

@@ 1 Introduction
Since their inception, literature reviews have been consistently used to streamline the advancement of various scientific fields (Snyder et al. 2016). Of these reviews, one of the most important types is the systematic literature review (SLR), which reproducibly synthesizes a significant portion of existing research relating to a specific research question being addressed (Kitchenham and Charters 2007; Bangdiwala 2024). This role has become increasingly more crucial, evidenced by how quite a few researchers consider them to be original research or potentially even a mandatory step in the scientific process itself (Kraus, Mahto, and Walsh 2023; Palmatier, Houston, and Hulland 2018).
Due to the large amount of new findings and research being disseminated through publications, it has become very difficult to release SLRs in a timely fashion (Qi, Li, and Lyu 2025; Ofori-Boateng et al. 2024; Tian et al. 2025). For example, the estimated time required to complete an SLR has increased significantly in the past few decades in the medical domain (Allen and Olkin 1999; Borah et al. 2017), which is further compounded by the necessity of using expert labor (Atkinson 2025). In tackling this, the vast majority of research relating to SLR automation focuses on the retrieval and screening of scientific papers (Orel et al. 2023; Marshall and Wallace 2019; Chen et al. 2025), as these are the most time-consuming steps (Chai et al. 2021). However, there remains the task of compiling the findings into a comprehensive review paper. Only a small number of works have been published (Kasanishi et al. 2023; Qi, Li, and Lyu 2025; Wang et al. 2024), let alone those which focus on the readability and hallucination mitigation aspects.
In this work, we present LiRA (Literature Review Agents), an agentic solution aimed at addressing the minimal research related to automated literature review writing. It is a large language model (LLM)-based agentic workflow building upon existing relevant works and integrates some of the most recent SLR-writing guidelines to generate accurate and high-quality reviews, with an additional emphasis on readability. Moreover, it is entirely out-of-the-box, requiring no task-specific pre-training or fine-tuning. We also reduce hallucination in the outputs, which is one of the main barriers in trustworthy automated writing and a key factor preventing the widespread use of similar artificial intelligence (AI)-powered systems (Alkaissi and McFarlane 2023; O’Connor et al. 2024; Xu et al. 2023).
@@ 2 Related Work
Comparisons have been made between agentic LLM systems and human cognition, as the idea of breaking down a task into smaller steps, which is often done by these workflows, is commonly used to describe how humans solve more intricate problems (Flower and Hayes 1981; Correa et al. 2023). Using this concept, several works show promise in implementing agentic workflows in various fields, such as the medical sciences (Tang et al. 2024) and law (Watson et al. 2025). In some cases, this results in an improvement of more than 90% compared to a simple baseline (Watson et al. 2025).
However, for the task of automated writing specifically, only few works have been published thus far (Qi, Li, and Lyu 2025; Wang et al. 2024; Shao et al. 2024; Tian et al. 2025). Of these works, only two of them address literature reviews or a similar document type and include open-source code (Qi, Li, and Lyu 2025; Wang et al. 2024). Neither paper takes output lengths and how they relate to the readability and evaluation of each proposed system using the listed metrics into account. As a result, no works seem to exist which tackle the issues of readability, and only minimal work exists in addressing the factuality of the resulting articles.
@@
For centuries (Lind 2014), the process of writing literature reviews has been considered crucial for the development of science (Meerpohl et al. 2012; Higgins et al. 2011; Chalmers, Hedges, and Cooper 2002), as it provides researchers valuable insight on which research questions to answer via an analysis of earlier works (Chalmers and Glasziou 2009; Eagly and Wood 1994). Moreover, improvements have been made to eliminate personal biases (Egger, Smith, and O’Rourke 2001) through the introduction of the SLR, which uses systematic methods of review for the collation and synthesis of findings (Randles and Finnegan 2023; Snyder 2019; Page et al. 2021).
Given how time-consuming (Borah et al. 2017) and costly (Michelson and Reuter 2019) this process is, a need for a viable alternative has emerged. Despite this, there is not much relevant innovation in natural language processing that tackles this issue (Mohammad et al. 2009; Kasanishi et al. 2023; Agarwal et al. 2011), let alone results indicating real-world usability. Therefore, we address all aforementioned problems by introducing an agentic workflow which uses LLMs to generate literature reviews automatically, demonstrating its potential by focusing on both readability and factuality.
@@ 3.1 Outline Drafter Agent
To manage the context size and maintain focus on the most relevant content, the outline is constructed from up to 50 references. We use existing heuristics (Wang et al. 2024) that recommend generating approximately 8 sections with around 4 subsections, but the agent can adapt this amount as needed. We also include default sections such as an introduction and conclusion to ensure consistency and completeness.
@@ 3.2 Subsection Writer Agent
Relevant references are retrieved using the FAISS index (Douze et al. 2025) based on section-level descriptions, capped at 25% of the total pool per subsection (min 3, max 150). This balances citation diversity with input tractability. Moreover, the reference titles, abstracts, or full texts can be retrieved depending on the availability. The agent outputs 1̃,000 words per subsection, enabling long-form synthesis while staying within model context limits. Also, the article title, abstract, and conclusion are written after the body is generated, mimicking human writing practices and avoiding the premature commitment of top-down generation.
@@ 3.4 Reviewer Agent
To improve factual accuracy and review quality, we introduce an LLM-based reviewer agent inspired by human editorial workflows. This agent evaluates intermediate outputs (e.g., outlines) based on adapted criteria from systematic review guidelines (Snyder 2019), including content completeness, transparency, clarity, and contribution.
@@ 3.6 Additional Implementation Details
All agents in LiRA were implemented using LangGraph. Moreover, each agent has its own memory, a standard practice for LLM-based agentic workflows to better act upon feedback (Sumers et al. 2023; Qian et al. 2024). This system of storing feedback in memory is comparable to Reflexion (Shinn et al. 2023), where the agent has to adjust its behavior based on the feedback provided. Parallelization is also included to increase the processing speed of certain steps, namely the researcher when analyzing papers and the content writer for generating the article (sub)sections.
Aspects from the design of MetaGPT (Hong et al. 2023) were adapted for more efficient inter-agent communication. Specifically, all agents are required to return their outputs as structured documents to avoid potential inefficiencies relating to information presentation, which are then sent to unique shared message pools for quick information retrieval. Moreover, the input contents are filtered based on the model’s maximum context window length (128,000 tokens in our case) to prevent information overload. As a method of improving the model’s output quality with minimal intervention, Zero-Shot Chain-of-Thought (Kojima et al. 2022) is included in the prompts for all agents except the researcher and editor, as they do not perform refinement.
@@ 4.1 Baselines
Introduced in an automated survey-writing challenge (Tian et al. 2025), MASS is the only agentic framework with publicly available code. Its workflow differs fundamentally from LiRA: Instead of decomposing the writing process into specialized roles with iterative feedback, MASS follows a strictly sequential pipeline. The system first clusters references by topical similarity to construct an outline, then generates section contents and a title directly from these clusters, and finally appends a conclusion. Long reference lists are handled by passing them as attachments.
@@
AutoSurvey (Wang et al. 2024) is a multi-stage framework for automatically generating literature surveys in computer science. Given a query (in our case, the original review title), it retrieves relevant publications and uses their titles and abstracts to construct an outline, followed by subsection drafting with refinement steps and partial use of retrieved article content (up to 1,500 tokens). While the paper claims retrieval-grounded drafting, we did not find corresponding functionality in the released code.
@@ 4.2 Metrics
We evaluate writing quality using both automatic and human assessments. For automatic evaluation, we use the Prometheus 2 LLM (Kim et al. 2024). It is an open-source LLM evaluator that uses the appropriate reference materials (the instruction, reference answer, and score rubric) to provide assessments which mostly align with those of human annotators. Three aspects are evaluated for all generated articles, namely the coverage, structure, and relevance. Here, coverage represents how broad the subject matter of the review is, while structure measures the organization and flow of the review, and relevance indicates how well the review is able to stay on-topic overall. Additional specifications for the model and how it measures the aforementioned aspects can be found in Appendix B.
@@ 4.3 Datasets
We evaluate the effectiveness of the LiRA framework using two datasets. The primary dataset is SciReviewGen (Kasanishi et al. 2023), a publicly available benchmark built on the Semantic Scholar Open Research Corpus (S2ORC) (Lo et al. 2020). It contains 10,000 review articles in computer science, referencing approximately 690,000 papers. Each review is annotated with structured metadata, including titles, section headers, full texts, and references. Following the setup in Shao et al. (2024), we randomly sample 125 reviews, ensuring each selected article contains a sufficient number of references—averaging around 70 per paper.
@@ 5.1 Using a different reviewer model
Based on concerns stemming from self-bias amplification (Xu et al. 2024), experimentation was performed using a reviewer model type different from the one used during generation for each component, which is based on existing suggestions. To this end, the gemma3:4b model was used through ollama (Team et al. 2025). This model was chosen because it is open-source, has a context window similar to gpt-4o-mini, and has lower hardware requirements compared to most other models. In addition, ollama was the selected model provider because of its accessibility and ease of use, with LangGraph already being compatible with it. The parameter values used by gemma3:4b were the default ollama ones aside from the context window size (128,000) and seed (42).
@@ Results
We examined if the results when using retrieval differed significantly compared to the baseline researcher setting, which was tested using the appropriate statistical tests. From the results (shown in 5), we note that only two results were significantly different compared to the baseline, indicating that LiRA is capable of performing similarly despite the different references used. More details on the statistical tests can be found in Appendix F.
@@ 6 Deployment
The deployment of LiRA will use the following steps. First, it will be developed using the Python version of LangGraph, which is an open-source and production ready agentic framework. Furthermore, the gpt-4o-mini LLM from AzureOpenAI will be used, with the possibility of using other models given LangGraph’s extensive support for various other endpoints such as form ollama.
@@ 7 Conclusion, Limitations, and Future Work
Several improvements could be made, mainly regarding the irreproducibility of results due to the usage of gpt-4o-mini for all experiments. This can be solved by using seedable models instead, which should be feasible given cunrrent LLM availability. In addition, there is a lack of open-source datasets for this task specifically, which hinders the generalizability of all results to other scientific fields. Therefore, we encourage authors to create additional datasets, ideally in a similar format to SciReviewGen, to facilitate the evaluation of similar systems in the future.
Furthermore, opportunities exist to create more end-to-end pipelines, as the current project does not take into account factors such as primary studies and risk of bias in randomized trials (i.e., the implementation of automated tools based on (Cochrane 2024)). Doing this would allow for the integration of more steps within the literature review writing process, namely the screening and search criteria definition steps, which would allow for better paper reproducibility.
@@ Acknowledgments
We would like to thank Marcela Haldan and Alexandra Noti for their help in providing annotations for the ScienceDirect articles.
@@ Appendix A Metrics Measuring the Similarity with Human-Written Articles
  • ROUGE-L: This metric is frequently used for text evaluation and functions by counting the number of recalled units in a text from the reference (Lin 2004). For this project, we used the ROUGE-L metric from the HuggingFace evaluate package.
  • Heading Soft Recall: Inspired by the idea of soft recall (Fränti and Mariescu-Istodor 2023), this metric quantitatively measures the heading coverage of a review without relying on exact matches, where a higher score indicates a more comprehensive outline by virtue of being similar to the human-written one. Here, we use the original implementation provided by Shao et al. (2024).
@@
  • Heading and Article Entity Recall: These metrics indicate the proportion of named entities in the baseline gold review mentioned in the generated text, between headings or the full review texts, respectively. The implementation for this is also taken from Shao et al. (2024). However, due to the presence of more scientific terms, we instead use the SciSpacy named entity tagger for all settings, specifically using the en_core_sci_lg model, as it has the largest available vocabulary.
@@ Appendix B LLM Evaluation
Here, we outline the details for LLM evaluation. For the specific LLM model, we used prometheus-7b-v2.0 using (Kim et al. 2024) with the same configurations outlined in Shao et al. (2024), namely a temperature of 0.01, cumulative probability of top tokens of 0.95, maximum number of new tokens of 512, and repetition penalty of 1.03. Also, because of hardware limitations, the model had to be run with fp8-quantization and a memory usage value of 0.8. Note that no seeding was used here, with the justification being that it would encourage more diverse outputs from the model, thus better mimicking reviews made by separate experts. Moreover, to simulate the process of evaluation by Subject Matter Experts (SMEs) and account for randomness, the evaluation was performed three times for all articles, with the final score being the average between all runs. Furthermore, due to the model’s context size, the evaluation for the criteria had to be done in the following ways:
@@
The rubric provided (shown in table 7) to the LLM contains scores ranging from 1 to 5, with the coverage, structure, and relevance evaluated for all generated literature reviews. Furthermore, the description for each score bracket is specified to accurately reflect the proper gradings. These were adapted from the rubric used by Wang et al. (2024), with minor adjustments to accommodate for the output structure. Also, because three runs were performed per setting, the results reported are the means of the average scores and the averaged resulting standard deviations.

### v2

@@ 1 Introduction
Since their inception, literature reviews have been consistently used to streamline the advancement of various scientific fields (snyder_identifying_2016). Of these reviews, one of the most important types is the Systematic Literature Review (SLR), which reproducibly synthesizes a significant portion of existing research relating to a specific research question being addressed (kitchenham_guidelines_2007; bangdiwala_importance_2024). This role has become increasingly more crucial, evidenced by how quite a few researchers consider them to be original research or potentially even a mandatory step in the scientific process itself (kraus_importance_2023; palmatier_review_2018).
Due to the large amount of new findings and research being disseminated through publications, it has become very difficult to release SLRs in a timely fashion (qi_generation_2025; ofori-boateng_towards_2024; tian_overview_2025). For example, the estimated time required to complete an SLR has increased significantly in the past few decades in the medical domain (allen_estimating_1999; borah_analysis_2017), which is further compounded by the necessity of using expert labor (atkinson_ai-pocalypse_2025). In tackling this, the vast majority of research relating to SLR automation focuses on the retrieval and screening of scientific papers (orel_automated_2023; marshall_toward_2019; chen_can_2025), as these are the most time-consuming steps (chai_research_2021). However, there remains the task of compiling the findings into a comprehensive review paper. Only a small number of works have been published (kasanishi_scireviewgen_2023; qi_generation_2025; wang_autosurvey_2024), let alone those which focus on the readability and hallucination mitigation aspects.
In this work, we present LiRA (Literature Review Agents), an agentic solution aimed at addressing the minimal research related to automated literature review writing. It is a Large Language Model (LLM)-based agentic workflow building upon existing relevant works and integrates some of the most recent SLR-writing guidelines to generate accurate and high-quality reviews, with an additional emphasis on readability. Moreover, it is entirely out-of-the-box, requiring no task-specific pre-training or fine-tuning. We also reduce hallucination in the outputs, which is one of the main barriers in trustworthy automated writing and a key factor preventing the widespread use of similar Artificial Intelligence (AI)-powered systems (alkaissi_artificial_2023; oconnor_large_2024; xu_critical_2023).
@@ 2 Related Work
Comparisons have been made between agentic LLM systems and human cognition, as the idea of breaking down a task into smaller steps, which is often done by these workflows, is commonly used to describe how humans solve more intricate problems (flower_cognitive_1981; correa_humans_2023). Using this concept, several works show promise in implementing agentic workflows in various fields, such as the medical sciences (tang_medagents_2024) and law (watson_law_2025). In some cases, this results in an improvement of more than 90% compared to a simple baseline (watson_law_2025).
However, for the task of automated writing specifically, only few works have been published thus far (qi_generation_2025; wang_autosurvey_2024; shao_assisting_2024; tian_overview_2025). Of these works, only two of them address literature reviews or a similar document type and include open-source code (qi_generation_2025; wang_autosurvey_2024). Neither paper takes output lengths and how they relate to the readability and evaluation of each proposed system using the listed metrics into account. As a result, no works seem to exist which tackle the issues of readability, and only minimal work exists in addressing the factuality of the resulting articles.
@@
For centuries (lind_treatise_2014), the process of writing literature reviews has been considered crucial for the development of science (meerpohl_scientific_2012; higgins_cochrane_2011; chalmers_brief_2002), as it provides researchers valuable insight on which research questions to answer via an analysis of earlier works (chalmers_avoidable_2009; eagly_using_1994). Moreover, improvements have been made to eliminate personal biases (egger_introduction_2001) through the introduction of the SLR, which uses systematic methods of review for the collation and synthesis of findings (randles_guidelines_2023; snyder_literature_2019; page_prisma_2021).
Given how time-consuming (borah_analysis_2017) and costly (michelson_significant_2019) this process is, a need for a viable alternative has emerged. Despite this, there is not much relevant innovation in natural language processing that tackles this issue (mohammad_using_2009; kasanishi_scireviewgen_2023; agarwal_towards_2011), let alone results indicating real-world usability. Therefore, we address all aforementioned problems by introducing an agentic workflow which uses LLMs to generate literature reviews automatically, demonstrating its potential by focusing on both readability and factuality.
@@ 3.1 Outline Drafter Agent
To manage the context size and maintain focus on the most relevant content, the outline is constructed from up to 50 references. We use existing heuristics (wang_autosurvey_2024) that recommend generating approximately 8 sections with around 4 subsections, but the agent can adapt this amount as needed. We also include default sections such as an introduction and conclusion to ensure consistency and completeness.
@@ 3.2 Subsection Writer Agent
Relevant references are retrieved using the FAISS index (douze_faiss_2025) based on section-level descriptions, capped at 25% of the total pool per subsection (min 3, max 150). This balances citation diversity with input tractability. Moreover, the reference titles, abstracts, or full texts can be retrieved depending on the availability. The agent outputs 1̃,000 words per subsection, enabling long-form synthesis while staying within model context limits. Also, the article title, abstract, and conclusion are written after the body is generated, mimicking human writing practices and avoiding the premature commitment of top-down generation.
@@ 3.4 Reviewer Agent
To improve factual accuracy and review quality, we introduce an LLM-based reviewer agent inspired by human editorial workflows. This agent evaluates intermediate outputs (e.g., outlines) based on adapted criteria from systematic review guidelines (snyder_literature_2019), including content completeness, transparency, clarity, and contribution.
@@ 3.6 Additional Implementation Details
All agents in LiRA were implemented using LangGraph. Moreover, each agent has its own memory, a standard practice for LLM-based agentic workflows to better act upon feedback (sumers_cognitive_2023; qian_chatdev_2024). This system of storing feedback in memory is comparable to Reflexion (shinn_reflexion_2023), where the agent has to adjust its behavior based on the feedback provided. Parallelization is also included to increase the processing speed of certain steps, namely the researcher when analyzing papers and the content writer for generating the article (sub)sections.
Aspects from the design of MetaGPT (hong_metagpt_2023) were adapted for more efficient inter-agent communication. Specifically, all agents are required to return their outputs as structured documents to avoid potential inefficiencies relating to information presentation, which are then sent to unique shared message pools for quick information retrieval. Moreover, the input contents are filtered based on the model’s maximum context window length (128,000 tokens in our case) to prevent information overload. As a method of improving the model’s output quality with minimal intervention, Zero-Shot Chain-of-Thought (kojima_large_2022) is included in the prompts for all agents except the researcher and editor, as they do not perform refinement.
@@ 4.1 Baselines
Introduced in an automated survey-writing challenge (tian_overview_2025), MASS is the only agentic framework with publicly available code. Its workflow differs fundamentally from LiRA: Instead of decomposing the writing process into specialized roles with iterative feedback, MASS follows a strictly sequential pipeline. The system first clusters references by topical similarity to construct an outline, then generates section contents and a title directly from these clusters, and finally appends a conclusion. Long reference lists are handled by passing them as attachments.
@@
AutoSurvey (wang_autosurvey_2024) is a multi-stage framework for automatically generating literature surveys in computer science. Given a query (in our case, the original review title), it retrieves relevant publications and uses their titles and abstracts to construct an outline, followed by subsection drafting with refinement steps and partial use of retrieved article content (up to 1,500 tokens). While the paper claims retrieval-grounded drafting, we did not find corresponding functionality in the released code.
@@ 4.2 Metrics
We evaluate writing quality using both automatic and human assessments. For automatic evaluation, we use the Prometheus 2 LLM (kim_prometheus_2024). It is an open-source LLM evaluator that uses the appropriate reference materials (the instruction, reference answer, and score rubric) to provide assessments which mostly align with those of human annotators. Three aspects are evaluated for all generated articles, namely the coverage, structure, and relevance. Here, coverage represents how broad the subject matter of the review is, while structure measures the organization and flow of the review, and relevance indicates how well the review is able to stay on-topic overall. Additional specifications for the model and how it measures the aforementioned aspects can be found in Appendix B.
@@ 4.3 Datasets
We evaluate the effectiveness of the LiRA framework using two datasets. The primary dataset is SciReviewGen (kasanishi_scireviewgen_2023), a publicly available benchmark built on the Semantic Scholar Open Research Corpus (S2ORC) (lo_s2orc_2020). It contains 10,000 review articles in computer science, referencing approximately 690,000 papers. Each review is annotated with structured metadata, including titles, section headers, full texts, and references. Following the setup in shao_assisting_2024, we randomly sample 125 reviews, ensuring each selected article contains a sufficient number of references—averaging around 70 per paper.
@@ 5.1 Using a different reviewer model
Based on concerns stemming from self-bias amplification (xu_pride_2024), experimentation was performed using a reviewer model type different from the one used during generation for each component, which is based on existing suggestions. To this end, the gemma3:4b model was used through ollama (team_gemma_2025). This model was chosen because it is open-source, has a context window similar to gpt-4o-mini, and has lower hardware requirements compared to most other models. In addition, ollama was the selected model provider because of its accessibility and ease of use, with LangGraph already being compatible with it. The parameter values used by gemma3:4b were the default ollama ones aside from the context window size (128,000) and seed (42).
@@ Results
We examined if the results when using retrieval differed significantly compared to the baseline researcher setting, which was tested using the appropriate statistical tests. From the results (shown in Table 5), we note that only two results were significantly different compared to the baseline, indicating that LiRA is capable of performing similarly despite the different references used. More details on the statistical tests can be found in Appendix F.
@@ 6 Deployment
The deployment of LiRA will use the following steps. First, it will be developed using the Python version of LangGraph, which is an open-source and production-ready agentic framework. Furthermore, the gpt-4o-mini LLM from AzureOpenAI will be used, with the possibility of using other models given LangGraph’s extensive support for various other endpoints such as form ollama.
@@ 7 Conclusion, Limitations, and Future Work
Several improvements could be made, mainly regarding the irreproducibility of results due to the usage of gpt-4o-mini for all experiments. This can be solved by using seedable models instead, which should be feasible given current LLM availability. In addition, there is a lack of open-source datasets for this task specifically, which hinders the generalizability of all results to other scientific fields. Therefore, we encourage authors to create additional datasets, ideally in a similar format to SciReviewGen, to facilitate the evaluation of similar systems in the future.
Furthermore, opportunities exist to create more end-to-end pipelines, as the current project does not take into account factors such as primary studies and risk of bias in randomized trials (i.e., the implementation of automated tools based on cochrane_cochrane_2024). Doing this would allow for the integration of more steps within the literature review writing process, namely the screening and search criteria definition steps, which would allow for better paper reproducibility.
@@ Acknowledgments
We would like to thank Marcela Haldan and Alexandra Noti for their help in providing annotations for the ScienceDirect articles. In addition, this research was (partially) supported by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union under grant agreements No. 101070212 (FINDHR) and No. 101201510 (UNITE). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of their respective employers, funders and/or granting authorities.
@@ Appendix A Metrics Measuring the Similarity with Human-Written Articles
  • ROUGE-L: This metric is frequently used for text evaluation and functions by counting the number of recalled units in a text from the reference (lin_rouge_2004). For this project, we used the ROUGE-L metric from the HuggingFace evaluate package.
  • Heading Soft Recall: Inspired by the idea of soft recall (franti_soft_2023), this metric quantitatively measures the heading coverage of a review without relying on exact matches, where a higher score indicates a more comprehensive outline by virtue of being similar to the human-written one. Here, we use the original implementation provided by shao_assisting_2024.
@@
  • Heading and Article Entity Recall: These metrics indicate the proportion of named entities in the baseline gold review mentioned in the generated text, between headings or the full review texts, respectively. The implementation for this is also taken from shao_assisting_2024. However, due to the presence of more scientific terms, we instead use the SciSpacy named entity tagger for all settings, specifically using the en_core_sci_lg model, as it has the largest available vocabulary.
@@ Appendix B LLM Evaluation
Here, we outline the details for LLM evaluation. For the specific LLM model, we used prometheus-7b-v2.0 using (kim_prometheus_2024) with the same configurations outlined in shao_assisting_2024, namely a temperature of 0.01, cumulative probability of top tokens of 0.95, maximum number of new tokens of 512, and repetition penalty of 1.03. Also, because of hardware limitations, the model had to be run with fp8-quantization and a memory usage value of 0.8. Note that no seeding was used here, with the justification being that it would encourage more diverse outputs from the model, thus better mimicking reviews made by separate experts. Moreover, to simulate the process of evaluation by Subject Matter Experts (SMEs) and account for randomness, the evaluation was performed three times for all articles, with the final score being the average between all runs. Furthermore, due to the model’s context size, the evaluation for the criteria had to be done in the following ways:
@@
The rubric provided (shown in table 7) to the LLM contains scores ranging from 1 to 5, with the coverage, structure, and relevance evaluated for all generated literature reviews. Furthermore, the description for each score bracket is specified to accurately reflect the proper gradings. These were adapted from the rubric used by wang_autosurvey_2024, with minor adjustments to accommodate for the output structure. Also, because three runs were performed per setting, the results reported are the means of the average scores and the averaged resulting standard deviations.